OpenStudy (anonymous):

Given the equation y - 3= 1/2 (x+6) in point-slope form, identify the equation of the same line in slope-intercept form. x − 2y = −12 y= 1/2x y=1/2x+6 y=1/2x+9

OpenStudy (anonymous):

Last one?

OpenStudy (anonymous):

Distribute the 1/2 y-3=1/2x-3 Add 3 to both sides y=1/2x

OpenStudy (anonymous):

It was wrong.

OpenStudy (anonymous):

I'm sorry....this was the answer I have.

OpenStudy (anonymous):

Standard form is of the form "Ax + By = C". y - 3 = (1/2)x + 3 y - (1/2)x = 6 Eliminate the fraction by multiplying both sides of the equation by 2. 2y - x = 12 Switch the x and y terms (the order) and multiply both sides so that the first term is positive. x - 2y = -12 (answer A)

OpenStudy (accidentalaichan):

@ZeruthC, that is the point-slope =form. Your slope-intercept form is y = mx +b